OpenStudy (anonymous):

Find all c in the interval [pi,2pi] such that 2cosx=1

OpenStudy (tkhunny):

This is a general knowledge problem, not a calculation problem. On \([\pi,2\pi]\), the cosine is positive only on \([3\pi/2,2\pi]\), so don't look in Quadrant III. One of the values you should just memorize it \(x = \pi/3\), which is the same as \(x = \ 2\pi - \pi/3 = 5\pi/3\). What is the cosine of that?

OpenStudy (anonymous):

I'm sorry i'm completely lost, I have no idea what you're talking about. is there one answer to this question or multiple?

OpenStudy (tkhunny):

Are you acquainted with the Unit Circle and its important points to memorize? That is how you solve this problem. Here's one. http://www.snow.edu/jonathanb/Courses/Math1060/unit_circ_trig.html Ponder it deeply. Where is the cosine equal to 1/2?